OpenStudy (anonymous):

simplify if this is possible log ^7√10

OpenStudy (anonymous):

\[\log ^{7}\sqrt{10}\]=\[(\frac{ 1 }{ 2 })^{7}\]

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

1/7? maybe

OpenStudy (anonymous):

you read the question wrong

OpenStudy (anonymous):

log is by itself...^7√10